# PUL-SLAM: Path-Uncertainty Co-Optimization with Lightweight Stagnation Detection for Efficient Robotic Exploration

### v1

@@ 4.1 Experimental Settings
To validate our algorithm, we established a ROS[28]-based simulation platform on an Ubuntu 20.04 system. This platform utilizes the Gazebo simulator to replicate realistic physical scenarios. Experiments were conducted using the TurtleBot3-Burger robot, equipped with a 360-degree LiDAR (maximum ranging distance of 3.5 meters) and wheel odometry for motion tracking. As shown in Fig. 2, the training scenario is a rectangular room with dimensions of 15 meters by 3 meters, filled with cylindrical obstacles. The density of these obstacles increases gradually from left to right, creating a gradient of complexity within the environment. The robot model initiates its learning and exploration process from the left side of the room. In the testing phase, three additional testing scenarios were introduced. These testing scenarios cover a range of areas from 56 to 128 square meters, representing environments of varying scales and complexities. Among them, Env-2 and Env-3 are standard scenarios commonly used for validating Active SLAM algorithms, and some studies[21, 23] have already conducted experiments in these environments. Env-4 is a complex suite designed by us, containing multiple obstacles with a more intricate layout. It is intended to further verify the robot’s adaptability and robustness in unknown and complex environments. By conducting tests in these diverse scenarios, we can comprehensively evaluate the performance of the proposed decision-making algorithm under different environmental conditions. The main parameters used in the experiment are shown in Table 1.
@@
| Env | Method | Time(s) | Distance(m) | Map Completeness(%) |
| Env-2 | Frontier | 322.14 | 50.19 | 99.30 |
| RRT | 553.68 | 58.40 | 99.24 | |
| DA-SLAM | 288.84 | 51.88 | 100 | |
| Ours | 235.23 | 41.84 | 98.56 | |
| Env-3 | Frontier | 334.98 | 34.10 | 99.68 |
| RRT | 268.89 | 37.29 | 98.45 | |
| DA-SLAM | 199.59 | 36.12 | 89.23 | |
| Ours | 167.07 | 29.26 | 94.13 | |
| Env-4 | Frontier | 778.67 | 60.32 | 99.55 |
| RRT | 698.30 | 90.15 | 99.85 | |
| DA-SLAM | 383.13 | 69.32 | 98.46 | |
@@ 4.2 Testing Evaluation
To comprehensively evaluate the performance of the proposed algorithm, this study designed a systematic comparative experimental framework. In three simulation environments (Env-2, Env-3, and Env-4), the proposed algorithm was compared against Frontier-based[8], RRT-based[12] and DA-SLAM (DRL-based)[21]. Three core metrics were selected: exploration time (from algorithm initiation to automatic termination), robot traversal distance, and map coverage ratio at algorithm termination. Ten independent trials were conducted per scenario to mitigate stochastic effects, with arithmetic means adopted as baseline performance measures. Experimental results are aggregated in Table LABEL:table:result, while Fig. 3 illustrates the exploration paths and mapping outcomes from representative trials (closest to mean performance), and Fig. 4 presents the exploration progress dynamics of each algorithm across different environments.
The results in Table LABEL:table:result show that the proposed algorithm achieves substantial improvements in both exploration time and path length across all test environments. In the most challenging Env-4, the proposed method completes exploration 65, 60, and 29 faster than Frontier, RRT, and DA-SLAM, respectively, while achieving a 42 shorter path than RRT and a 25 reduction compared to DA-SLAM. Consistent gains are observed in Env-2 and Env-3, with 25–30 time savings and approximately 20 shorter paths. While map completeness is slightly lower than baselines in Env-2 and Env-3, this reflects the algorithm’s deliberate trade-off: by suppressing redundant revisits and prioritizing unexplored regions, it optimizes the balance between efficiency and coverage.
The exploration trajectory visualization in Fig. 3 intuitively demonstrates the path optimization advantages of the proposed algorithm. Compared to the baseline methods, the proposed algorithm’s exploration path exhibits a highly concentrated characteristic, effectively reducing backtracking and detouring phenomena. Specifically, the Frontier method shows evident local oscillations and redundant regional exploration during the process; RRT frequently generates tortuous paths, demonstrating significant path redundancy; DA-SLAM, while achieving higher coverage, fails to terminate the exploration process in a timely manner, resulting in lower path efficiency. In contrast, the proposed algorithm achieves efficient environmental coverage through smooth and direct exploration trajectories.
Fig. 4 provides comprehensive insights into the exploration progress dynamics of the four algorithms across varying environmental complexities. The left subplots (coverage versus time) show that the proposed method consistently attains high coverage faster than all baselines. In Env-2 and Env-3, it reaches near-complete coverage in 250 seconds and 175 seconds, respectively, substantially outpacing Frontier, RRT, and DA-SLAM. In the highly cluttered Env-4, it achieves nearly 100 coverage within 300 seconds, while the others require 400 to 800 seconds. The right subplots (coverage versus path length) further confirm its path efficiency: the proposed algorithm achieves comparable or higher coverage with significantly shorter trajectories—approximately 40 meters in Env-2, under 30 meters in Env-3, and only 50 meters in Env-4, compared to 50–70 meters for the baselines. This demonstrates its superior balance between rapid exploration and motion cost.
@@ 4.3 Ablation Experiment
We systematically validate the performance improvements brought by the Path-Uncertainty Co-Optimization Reward (PUR) and the LSD through ablative experiments. Three configurations were established:
@@
Baseline Method: Uncertainty-based DRL[21] (blue curve)
@@
LSD-Enhanced Method: Baseline + LSD (green curve)
@@
Full Method: PUR + LSD (red curve)
Fig. 5 depicts the reward curves during training. The LSD module significantly suppresses abnormal reward fluctuations, attributed to its laser-based static anomaly detection mechanism and map-update stagnation detection, which promptly trigger action correction to avoid learning ineffective exploration behaviors, thereby reducing reward oscillations. The PUR framework further enhances training efficiency by dynamically balancing map information gain and path penalties, enabling more efficient allocation of exploration resources.
Fig. 6 presents multiple performance metrics. Regarding exploration time, both the LSD-enhanced and PUR methods significantly outperform the baseline. Notably, the LSD method exhibits shorter exploration time than PUR. This apparent discrepancy is explained by analyzing map completeness (Fig. 6(b)) and movement distance (Fig. 6(c)): The PUR method achieves faster convergence, attaining higher map completeness within the same training episodes (the red curve in Fig. 6(b) reaches high coverage earlier). This thorough exploration behavior increases exploration time and movement distance (higher distance values for the red curve in Fig. 6(c)). In contrast, while the LSD method optimizes stagnation issues, its lack of dynamic reward guidance limits exploration scope within the same episodes, resulting in shorter time but incomplete map coverage. The baseline method performs worst across all metrics, exhibiting substantial time waste and redundant movement, underscoring the necessity of action correction and reward optimization.
In summary, the LSD and PUR modules enhance system performance at the execution layer and decision layer, respectively. Their synergy significantly improves training efficiency and final algorithmic performance.
@@ 4.4 Real-world Experiment
Fig. 7 presents the records from the real-world experiment. Fig. 7(a) shows the laboratory environment, clearly displaying the densely distributed obstacles and complex spatial structures; Fig. 7(b) illustrates the experimental platform; and Fig. 7(c) presents the robot’s final exploration trajectory (green path) along with the constructed complete environmental map. The experimental results indicate that the robot completed the entire environment exploration within 122.32 seconds, with a final path length of 15.96 meters, while successfully avoiding collisions with various obstacles in the environment. This result shows high consistency with the performance observed in the simulation environment, validating the transferability of the proposed algorithm from simulation to real-world settings. Notably, the robot’s exploration path during the process exhibited a highly concentrated characteristic, significantly reducing unnecessary backtracking and detouring behaviors.
@@ 5 CONCLUSIONS
This paper proposes a Path-Uncertainty Co-Optimization DRL framework with Lightweight Stagnation Detection mechanism, which significantly enhances robotic exploration efficiency. Experimental results show that the proposed method substantially outperforms the frontier-based method, the RRT-based approach and DA-SLAM in terms of exploration efficiency, while maintaining reliable map completeness. Specifically, it reduces exploration time by up to 65 relative to the frontier-based method and shortens path length by up to 42 compared to RRT, consistently outperforming all baselines across diverse and complex scenarios. Ablation studies further confirm the complementary roles of the Path-Uncertainty Co-Optimization Reward and the Lightweight Stagnation Detection mechanism: PCR enhances the efficient allocation of exploration resources at the decision-making level, while LSD suppresses ineffective behaviors at the execution level. Their synergistic integration significantly accelerates training convergence. Moreover, real-world experiments on a physical robotic platform validate the successful sim-to-real transferability of the proposed approach, demonstrating its practical applicability and deployment potential. Future work will focus on further optimizing coverage uniformity and enhancing system robustness in dynamic and large-scale environments.

### v2

@@ 4.1 Experimental Settings
To validate our algorithm, we established a ROS[quigley2009ros]-based simulation platform on an Ubuntu 20.04 system. This platform utilizes the Gazebo simulator to replicate realistic physical scenarios. Experiments were conducted using the TurtleBot3-Burger robot, equipped with a 360-degree LiDAR (maximum ranging distance of 3.5 meters) and wheel odometry for motion tracking. As shown in Fig. 3, the training scenario is a rectangular room with dimensions of 15 meters by 3 meters, filled with cylindrical obstacles. The density of these obstacles increases gradually from left to right, creating a gradient of complexity within the environment. The robot model initiates its learning and exploration process from the left side of the room. In the testing phase, three additional testing scenarios were introduced. These testing scenarios cover a range of areas from 56 to 128 square meters, representing environments of varying scales and complexities. Among them, Env-2 and Env-3 are standard scenarios commonly used for validating Active SLAM algorithms, and some studies[alcalde2022slam, placed2020deep] have already conducted experiments in these environments. Env-4 is a complex suite designed by us, containing multiple obstacles with a more intricate layout. It is intended to further verify the robot’s adaptability and robustness in unknown and complex environments. By conducting tests in these diverse scenarios, we can comprehensively evaluate the performance of the proposed decision-making algorithm under different environmental conditions. The main parameters used in the experiment are shown in Table 1.
@@
| Env | Scenario | Method | Time(s) | Path Length(m) | Map Completeness (%) |
| Env-2 |
Frontier | 322.14 | 50.19 | 99.30 |
| RRT | 553.68 | 58.40 | 99.24 | ||
| DA-SLAM | 288.84 | 51.88 | 100 | ||
| Ours | 235.23 | 41.84 | 98.56 | ||
| Env-3 |
Frontier | 334.98 | 34.10 | 99.68 |
| RRT | 268.89 | 37.29 | 98.45 | ||
| DA-SLAM | 199.59 | 36.12 | 89.23 | ||
| Ours | 167.07 | 29.26 | 94.13 | ||
| Env-4 |
Frontier | 778.67 | 60.32 | 99.55 |
| RRT | 698.30 | 90.15 | 99.85 | ||
| DA-SLAM | 383.13 | 69.32 | 98.46 | ||
@@ 4.2 Testing Evaluation
To comprehensively evaluate the performance of the proposed algorithm, this study designed a systematic comparative experimental framework. In three simulation environments (Env-2, Env-3, and Env-4), the proposed algorithm was compared against Frontier-based[yamauchi1997frontier], RRT-based[umari2017autonomous] and DA-SLAM (DRL-based)[alcalde2022slam]. Three core metrics were selected: exploration time (from algorithm initiation to automatic termination), robot traversal distance, and map coverage ratio at algorithm termination. Ten independent trials were conducted per scenario to mitigate stochastic effects, with arithmetic means adopted as baseline performance measures. Experimental results are aggregated in Table 2, while Figs. 4, 5 and 6 illustrates the exploration paths and mapping outcomes from representative trials (closest to mean performance), and Fig. 7 presents the exploration progress dynamics of each algorithm across different environments.
The results in Table 2 show that the proposed algorithm achieves substantial improvements in both exploration time and path length across all test environments. In the most challenging Env-4, the proposed method completes exploration 65, 60, and 29 faster than Frontier, RRT, and DA-SLAM, respectively, while achieving a 42 shorter path than RRT and a 25 reduction compared to DA-SLAM. Consistent gains are observed in Env-2 and Env-3, with 25–30 time savings and approximately 20 shorter paths. While map completeness is slightly lower than baselines in Env-2 and Env-3, this reflects the algorithm’s deliberate trade-off: by suppressing redundant revisits and prioritizing unexplored regions, it optimizes the balance between efficiency and coverage.
@@ 4.3 Ablation Experiment
We conducted systematic ablation studies to quantitatively evaluate the contribution of each component in our proposed framework. The experiment was designed with three distinct configurations to isolate the effects of our key innovations:
@@
Baseline Method: Uncertainty-based DRL approach [alcalde2022slam] (blue curves)
@@
LSD-Enhanced Method: Baseline + Lightweight Stagnation Detection (green curve)
@@
Full Method: Path-Uncertainty Co-Optimization Reward (PUR) + Lightweight Stagnation Detection (red curve)
Figure 8 presents the average episode rewards over 2000 training episodes. The baseline method demonstrates extremely poor convergence characteristics throughout the entire training process. Despite 2000 episodes of training, its reward curve continues to exhibit severe oscillations, failing to achieve a stable policy. This instability stems from the algorithm’s inability to recognize and correct inefficient exploration patterns, allowing the agent to repeatedly learn destructive policies that result in environmental collisions or exploration stagnation. In contrast, both LSD-enhanced variants exhibit significantly stabilized training curves. This stability is attributed to the dual detection mechanism of the LSD module: LiDAR Static Anomaly Detection promptly identifies physical motion failures (such as wheel slippage or collisions), while Map Update Stagnation Detection terminates episodes where exploration progress stagnates despite active movement commands. By terminating these episodes characterized by motion failures or map update stagnation, the mechanism prevents the reinforcement learning agent from incorporating these detrimental experiences into its policy updates. The Full Method not only demonstrates superior stability but also achieves accelerated convergence. The path penalty term (Eq. 2) actively discourages inefficient exploration behaviors, while the D-optimality criterion ensures sufficient uncertainty reduction for reliable mapping.
Fig. 9 presents a comprehensive analysis of our ablation study across three critical performance metrics. Regarding exploration time, both the LSD-enhanced and the Full methods significantly outperform the baseline. Notably, the LSD-enhanced method exhibits shorter exploration time than the Full method. This apparent discrepancy is explained by analyzing map completeness (Fig. 9(b)) and movement distance (Fig. 9(c)): The Full method achieves faster convergence, attaining higher map completeness within the same training episodes (the red curve in Fig. 9(b) reaches high coverage earlier). This thorough exploration behavior increases exploration time and movement distance (higher distance values for the red curve in Fig. 9(c)). In contrast, while the LSD-enhanced method optimizes stagnation issues, its lack of dynamic reward guidance limits exploration scope within the same episodes, resulting in shorter time but incomplete map coverage. The baseline method performs worst across all metrics, exhibiting substantial time waste and redundant movement, underscoring the necessity of action correction and reward optimization.
This analysis validates the synergistic relationship between our two proposed mechanisms: the LSD module provides essential execution-level stability by filtering out catastrophic failures, while the DRL framework optimizes decision-making efficiency through a balanced reward structure. Their integration creates a learning environment where policy updates are derived from consistently productive exploration experiences, thereby ensuring training stability and accelerating convergence speed.
@@ 4.4 Real-world Experiment
Fig. 10 shows the laboratory environment, clearly displaying the densely distributed obstacles and complex spatial structures. Fig. 11 illustrates the experimental platform. Fig. 12 presents the robot’s final exploration trajectory (green path) along with the constructed complete environmental map. The experimental results indicate that the robot completed the entire environment exploration within 122.32 seconds, with a final path length of 15.96 meters, while successfully avoiding collisions with various obstacles in the environment. This result shows high consistency with the performance observed in the simulation environment, validating the transferability of the proposed algorithm from simulation to real-world settings. Notably, the robot’s exploration path during the process exhibited a highly concentrated characteristic, significantly reducing unnecessary backtracking and detouring behaviors.
@@ 5 CONCLUSIONS
This paper proposes a Path-Uncertainty Co-Optimization DRL framework with Lightweight Stagnation Detection mechanism, which significantly enhances robotic exploration efficiency. Experimental results show that the proposed method substantially outperforms the frontier-based method, the RRT-based approach and DA-SLAM in terms of exploration efficiency, while maintaining reliable map completeness. Specifically, it reduces exploration time by up to 65 relative to the frontier-based method and shortens path length by up to 42 compared to RRT, consistently outperforming all baselines across diverse and complex scenarios. Ablation studies further confirm the complementary roles of the Path-Uncertainty Co-Optimization Reward and the Lightweight Stagnation Detection mechanism: PUR enhances the efficient allocation of exploration resources at the decision-making level, while LSD suppresses ineffective behaviors at the execution level. Their synergistic integration significantly accelerates training convergence. Moreover, real-world experiments on a physical robotic platform validate the successful sim-to-real transferability of the proposed approach, demonstrating its practical applicability and deployment potential. Future work will focus on further optimizing coverage uniformity and enhancing system robustness in large-scale environments.